Determine the intercepts for the linear rule, 4y - 7x= 28
Answer by checkley71(8403)

4Y-7X=28 OR IN THE FORM OF Y=mX+b WE HAVE 4Y=7X+28 OR Y=7/4*X+7 THUS THE Y INTERCEPT IS 7 AND SETTING Y=0 WE GWT 0=7/4*X+7 OR 7/4*X=-7 OR X=-7(4/7) OR
X=-4 IS THE X INTERCEP